# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Medical Image

### v1

@@ 6 User Study
Sec. per task ($\downarrow$) $131.7\pm 168.5$ $94.2\pm 89.0$
No. iterations ($\downarrow$) $6(47)$ $3(19)$
@@
Results. Participants produced more accurate segmentations (0.84 vs. 0.73 Dice; $p=0.001$ using a paired t-test) using ScribblePrompt (Table 2). Participants spent $\approx 1.5$ minutes per segmentation on average using ScribblePrompt, compared to over 2 minutes per segmentation with SAM ($p=0.02$ using a paired t-test). While using ScribblePrompt, participants updated the prediction fewer times before being satisfied with the segmentation.
@@
Upon completion, 15 out of 16 participants reported they preferred using ScribblePrompt to SAM and one participant had no preference. All participants reported it was easier to achieve the target segmentation using ScribblePrompt vs. SAM. 94.8% of participants reported ScribblePrompt was better than SAM at refining its predictions in response to scribbles, and 87.5% preferred using ScribblePrompt for clicks. 93.8% of participants reported that they felt it took less time to achieve the target segmentation using ScribblePrompt compared to using SAM.
@@
Discussion. The most common factors that influenced participants preference for ScribblePrompt were 1) being able use multiple types of prompts including scribbles and 2) ScribblePrompt’s responsiveness to their corrections, which enabled fine-grained control over the next prediction. For some tasks, such as retinal vein segmentation, SAM was unable to make accurate predictions even with many corrections. As a result, there was more variability in segmentation accuracy and time spent per image, for SAM than ScribblePrompt.
7 Inference Runtime
@@
Setup. We evaluate computational efficiency by measuring inference time on a single CPU for one prediction with a scribble input covering 128 pixels. Performance on a CPU reflects practical utility: because of a variety of barriers such as protected health information, users may not be able to send their data to a server for computation and must rely on local, potentially CPU-only, computing.
@@
Results. We report performance results in Table 3. On a single CPU, ScribblePrompt-Unet requires $0.27$ seconds on average per prediction, enabling the model to be used in real-time even in low-resource environments. The next most accurate model on simulated scribbles, SAM ViT-h, requires over 2 minutes on average per prediction on a CPU. We show results with similar trends for GPU hardware in Appendix E.
@@
ScribblePrompt-Unet achieves low latency inference because of its efficient fully-convolutional architecture. With SAM and SAM-Med2d, latency scales with the number of interactions because each point is encoded as a 256-dimensional vector embedding. For ScribblePrompt-Unet and MIDeepSeg, clicks and scribbles are encoded in masks, so inference time is constant with the number of interactions.
Table 3: Performance Summary. We measure inference time on a single CPU for a prediction with a random scribble input covering 128 pixels. We report mean and standard deviation across 1,000 runs. ScribblePrompt-SAM has the same runtime and number of parameters as SAM ViT-b.
Model Param. Runtime (sec)
SAM (ViT-h) 641M $130.79\pm 7.96$
SAM (ViT-b) 94M $13.59\pm 0.77$
SAM-Med2d w/ adapter 271M $1.23\pm 0.07$
SAM-Med2d w.o. adapter 91M $0.63\pm 0.02$
MIDeepSeg 3M $0.08\pm 0.02$
ScribblePrompt Unet 4M $0.27\pm 0.04$
@@ 8 Ablations
We conduct two ablations of ScribblePrompt: (1) varying the percentage of synthetic label inputs used during training, and (2) varying the types of prompts simulated during training. In these ablations, we report results on the validation data, specifically the validation splits of the nine validation datasets that were not used in training.
@@ 8.1 Synthetic Labels
To improve generalization to unseen tasks and datasets, we introduced synthetic labels during training. Training with some synthetic labels helps to prevent the model from over fitting, especially for small single-label datasets.
Setup. We trained ScribblePrompt-Unet with probability $p_{synth}\in\{0,0.25,0.5,1.0\}$. We evaluate models using Dice after five steps following the same six simulated interaction procedures in the main results. We show results for ScribblePrompt-SAM in Appendix B.3.2.
Results. Figure 5 shows the effect of varying the percentage of synthetic labels during training for ScribblePrompt-Unet. Using $p_{synth}=0.5$ results in the highest Dice on validation data for all simulated interaction procedures.
Figure 5: Training with some synthetic labels improves Dice. Mean change in weighted Dice relative to ScribblePrompt-Unet trained without any synthetic labels ($p_{synth}=0$). We show results for six different interaction simulation procedures at inference-time. Error bars show 95% CI. Training with both real and synthetic labels improves generalization to new datasets, compared to training with only real labels. Training with only synthetic labels results in worse Dice on new datasets compared to training with real labels.
@@ 8.2 Prompt Types
Setup. We evaluate ScribblePrompt-Unet models trained with different combinations of prompts. We consider three models trained on different subsets of prompt types:
  • ScribblePrompt-Unet (scribbles) was trained on bounding boxes and scribbles.
  • ScribblePrompt-Unet (clicks) was trained on bounding boxes and clicks following our click simulation algorithms.
  • ScribblePrompt-Unet (random clicks) was trained on bounding boxes and only random clicks.
We compare these variants of ScribblePrompt-Unet to our final ScribblePrompt-Unet model, trained on scribbles, clicks and bounding boxes.
Results. Figure 6 shows that our final version of ScribblePrompt-Unet trained with scribbles, clicks and bounding boxes predicts segmentations more accurately than do versions of ScribblePrompt-Unet trained with fewer prompt types. ScribblePrompt-Unet trained on scribbles (no clicks), has similar performance to ScribblePrompt-Unet on scribbles but worse performance on clicks. Versions of ScribblePrompt-Unet trained on our mixture of clicks or only random clicks have similar performance to ScribblePrompt-Unet on clicks.
Figure 6: Varying the types of simulated prompts during training. ScribblePrompt-Unet has similar or better performance than models trained on scribbles (no clicks) and models trained on clicks (no scribbles) for both click-focused and scribble-focused interaction scenarios. We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. We calculate mean Dice on validation data weighting each dataset equally. Errorbars show 95% CI.
@@ 9 Conclusion
We present ScribblePrompt, a framework for interactive segmentation that enables humans to segment diverse medical images with scribbles, clicks and bounding boxes. We introduce methods for simulating realistic user interactions and generating synthetic labels. These methods enable us to train ScribblePrompt models that generalize to unseen segmentation tasks and datasets. ScribblePrompt models are more accurate than baselines, and ScribblePrompt-Unet is very computationally efficient, even on a CPU. We show in a user study that the majority of users prefer ScribblePrompt and achieve segmentations with 15% higher Dice with less perceived effort than the next most accurate baseline. ScribblePrompt promises to significantly reduce the burden of manual segmentation in medical imaging.
@@ 10 Acknowledgements
This work was supported in part by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard as well as Quanta Computer Inc. and the Wistron Corporation. Research reported in this paper was supported by the National Institute of Biomedical Imaging and Bioengineering of the National Institutes of Health under award number R01EB033773. Much of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center.
@@
Appendix A Demo
We showcase ScribblePrompt in a demonstration video and interactive online demo available at scribbleprompt.csail.mit.edu. The video shows ScribblePrompt-Unet being used in real time to segment test images from datasets not seen by the model during training.
Appendix B Details About ScribblePrompt Method
@@ B.1.1 Scribbles
Line Scribbles. Figure 7 illustrates the process of simulating line scribbles.
Figure 7: Line scribbles. Given an input mask $z$, we draw random lines by sampling two end points from $\{(u,v)|z_{uv}=1\}$. We use a random deformation field to warp the line scribbles and then multiply by the binary input mask $z$ to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (top) and negative scribbles by applying the algorithm to the background $1-y$ (bottom).
Centerline Scribbles. Figure 8 illustrates the process of simulating centerline scribbles.
@@
Contour Scribbles. Figure 9 illustrates the process of simulating contour scribbles.
Figure 9: Contour scribbles. We simulate a rough contour of the desired segmentation within the boundaries of the label. Given a mask $z$, we first blur the reverse mask $\bar{z}$ to reduce the size of the label such that $\tilde{z}:=1-\max(\bar{z},\bar{z}\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{z}<h$ sampled in some intensity range $h\sim U[\tilde{z}_{min},\tilde{z}_{max}]$, and use an edge detection algorithm to extract a contour inside the boundary of the given mask $z$. We break up the contour using a random mask and use a random deformation field to warp the broken contour. Lastly, we multiply the scribble mask by the input binary mask to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (bottom) and negative scribbles by applying the algorithm to the background $1-y$ (top).
@@ B.1.2 Clicks
Interior Border Region Clicks. Figure 10 illustrates the process for simulating interior border region clicks.
Figure 10: Interior border region clicks. We sample clicks from a border region inside the boundary of a given mask. Given a mask $z$, we first blur the reverse mask $1-z$ to reduce the size of the label such that $\tilde{z}:=1-\max(1-z,(1-z)\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel with $\sigma\sim U[\sigma_{min},\sigma_{max}]$. We then sample click coordinates from $\{(u,v)|\tilde{z}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{z}_{min},\tilde{z}_{max})$. We show the simulation process for negative border region clicks on the background $1-y$ (top) and positive border region clicks on the label $y$ (bottom).
@@ B.2 Architecture and Training
Normalization Layers. In preliminary experiments, we evaluated normalization layers in the ScribblePrompt-Unet architecture such as Batch Norm [52], Instance Norm [124], Layer Norm [8], and Channel Norm [136]. Including normalization did not improve the mean Dice on validation data compared to using no normalization layers (Figure 11).
@@
Figure 11: Normalization Layers. Mean Dice recorded while training ScribblePrompt-Unet with different normalization layers, averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
@@
Loss Function. In preliminary experiments, we trained ScribblePrompt with Soft Dice Loss [32], a combination of Soft Dice Loss and Binary Cross-Entropy Loss, and a combination of Soft Dice Loss and Focal Loss [72], similar to [61]. In the latter two losses, Dice Loss and BCE Loss or Focal Loss are weighted equally. We found that the combination of Soft Dice Loss and Focal Loss resulted in slightly higher mean Dice on the validation data for ScribblePrompt-Unet and ScribblePrompt-SAM. Figure 12 shows Dice recorded during training in preliminary experiments with ScribblePrompt-Unet.
@@
Figure 12: Varying the Loss. Mean Dice recorded while training ScribblePrompt-Unet with different loss functions. We report Dice averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
@@
ScribblePrompt-SAM Details. To train ScribblePrompt-SAM, we took the pre-trained weights from the Segment Anything Model [61] with ViT-b backbone and froze all components of the network except for the decoder.
@@
The SAM architecture can make predictions in “single-mask mode” or “multi-mask mode.” In “single-mask mode,” the decoder outputs a single predicted segmentation given an input image and user interactions. In “multi-mask mode,” the decoder predicts three possible segmentations and then outputs the segmentation with the highest predicted IoU by a MLP. We trained and evaluated ScribblePrompt-SAM in “multi-mask mode”. During training we included a MSE term in the segmentation loss to train the MLP to predict the IoU of the predictions, as in [61].
@@ B.3 Synthetic Labels
To help reduce task overfitting – memorizing the segmentation task for single-label datasets and thus ignoring user prompts – we introduce a mechanism to generate synthetic labels. During training, for a given sample $(x^{t},y^{t})$, with probability $p_{synth}$ we replace $y^{t}$ with a synthetic label $y^{\prime}$.
B.3.1 Implementation
@@
We use a superpixel algorithm [33] with randomly sampled scale parameter $\lambda\sim U[1,500]$ to partition the image $x^{t}$ into a map of $k$ superpixels, $z\in\{1,\dots,k\}^{n\times n}$. Then, we randomly select a superpixel as the synthetic label $y^{\prime}:=\mathbbm{1}[z=k]$. Figure 14 shows examples of training images and the corresponding maps of possible synthetic labels with different $\lambda$.
B.3.2 Synthetic Labels with ScribblePrompt-SAM
Setup. We train models with different percentages of synthetic labels during training by varying $p_{synth}$. We trained ScribblePrompt-SAM with probability $p_{synth}\in\{0,0.25,0.5\}$. We evaluate models using Dice after five steps following the same six simulated interaction procedures in the main results.
Results. Training with some synthetic labels also improves ScribblePrompt-SAM’s performance on validation data from nine (validation) datasets not seen during training (Figure 13). For ScribblePrompt-SAM, training with 50% synthetic labels leads to the highest Dice on unseen datasets at inference time.
Figure 13: Varying the amount of synthetic labels during training for ScribblePrompt-SAM. We report change in Dice relative to ScribblePrompt-SAM trained without any synthetic labels ($p_{synth}=0$). We show Dice after five steps of simulated interactions following six different (inference-time) interaction procedures. Errorbars show 95% CI.

### v2

@@ 5.3 User Study
HD95 ($\downarrow$) $16.1\pm 37.1$ $4.5\pm 9.1$
Sec./task ($\downarrow$) $131.7\pm 168.5$ $94.2\pm 89.0$
Iter./task ($\downarrow$) $6~{}(47)$ $3~{}(19)$
To assess the practical utility of ScribblePrompt, we conducted a user study with experienced annotators. We compare ScribblePrompt-UNet to SAM (ViT-b). We selected ScribblePrompt-UNet for the user study because it had similar performance and lower latency compared to ScribblePrompt-SAM. We selected SAM (ViT-b) because it had the highest Dice score in our experiments with clicks and lower latency than the next closest baseline, SAM (ViT-h).
Setup. Study participants ($n=16$) were neuroimaging researchers at an academic hospital. They used each model to segment a series of nine test images from different evaluation datasets. For each segmentation task, the participants were shown the target segmentation and were asked to prompt the model until the predicted segmentation closely matched the target, or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation. We provide additional details and visualizations in Appendix G.
@@
Results. Participants produced more accurate segmentations (0.84 vs. 0.73 Dice; $p=0.001$ using a paired t-test) using ScribblePrompt (Tab. 2). Participants spent $\approx 1.5$ minutes per segmentation on average using ScribblePrompt, compared to over 2 minutes per segmentation with SAM ($p=0.02$ using a paired t-test). While using ScribblePrompt, participants updated the prediction fewer times before being satisfied with the segmentation.
@@
Upon completion, 15 out of 16 participants reported they preferred using ScribblePrompt to SAM and one participant had no preference. All participants reported it was easier to achieve the target segmentation using ScribblePrompt compared to SAM. 93.8% of participants reported ScribblePrompt was better than SAM at refining its predictions in response to scribbles. 87.5% of participants preferred using ScribblePrompt over SAM for clicks.
@@
Discussion. The most common factors that influenced participant preference for ScribblePrompt were 1) being able get accurate predictions from multiple types of prompts, including scribbles and 2) responsiveness to their corrections, which enabled fine-grained control over the next prediction. For some tasks, such as retinal vein segmentation, SAM was unable to make accurate predictions even with many corrections (Fig. 35). As a result, there was more variability in Dice score and time per segmentation, for SAM than ScribblePrompt.
5.4 Inference Runtime
@@
Setup. We evaluate computational efficiency by measuring inference time on a single CPU for one prediction with a scribble input covering 128 pixels. Performance on a CPU reflects practical utility. Because of a variety of barriers such as protected health information, users may not be able to send their data to a server for computation, and must rely on local, most often CPU-only, computing.
@@
Results. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used in real-time even in low-resource environments. The next most accurate model (Fig. 6), SAM ViT-h, requires over 2 minutes per prediction on a CPU ($130.79\pm 7.96$ s). SAM ViT-b takes $13.59\pm 0.77$ seconds per prediction. We report results for all baselines and on GPU hardware with similar trends in Appendix H.
@@
Its efficient fully-convolutional architecture gives ScribblePrompt-UNet low latency inference. With SAM, latency scales with the number of interactions because each point is encoded as a 256-dimensional vector embedding. For ScribblePrompt-UNet, clicks and scribbles are encoded in masks, so inference time (per prediction) is constant with the number of interactions.
(a) Probability of synthetic labels.
(b) Interactions during training.
Figure 7: Ablations. We report Dice after five steps of simulated interactions following three inference-time interaction procedures. Error bars show 95% CI. (a) shows mean change in Dice relative to ScribblePrompt-UNet trained without any synthetic labels ($p_{synth}=0$).
@@ 6 Ablations
We conduct two ablations of important design decisions: (1) synthetic label inputs used during training, and (2) types of prompts simulated during training. We report results on the validation splits of nine datasets unseen during training.
@@ 6.1 Synthetic Labels
Setup. We trained ScribblePrompt-UNet with different values of $p_{synth}$, the probability of sampling a synthetic label.
Results. Fig. 7(a) shows the effect of varying the percentage of synthetic labels during training for ScribblePrompt-UNet. Training with both real and synthetic labels improves generalization to new datasets, compared to training with only real labels. Using $p_{synth}=0.5$ results in the highest Dice. Training with only synthetic labels results in worse Dice scores. We show similar results for ScribblePrompt-SAM and other interaction procedures in Sec. I.1.
@@ 6.2 Prompt Types
Setup. We evaluate ScribblePrompt-UNet models trained with different combinations of prompts, compared to the complete ScribblePrompt-UNet:
  • ScribblePrompt-UNet (scribbles) trained on boxes and scribbles.
  • ScribblePrompt-UNet (clicks) trained on boxes and clicks.
  • ScribblePrompt-UNet (random clicks) trained on boxes and random clicks.
Results. Fig. 7(b) shows ScribblePrompt-UNet trained with scribbles, clicks and bounding boxes predicts segmentations more accurately than do ablated versions of ScribblePrompt-UNet. We show results for other interaction procedures with similar trends in Sec. I.2.
@@ 7 Conclusion
We present ScribblePrompt, a practical framework for interactive segmentation that enables users to segment diverse medical images with scribbles, clicks, and bounding boxes. We introduce methods for simulating realistic user interactions and generating synthetic labels. These methods enable us to train models that generalize to unseen segmentation tasks and datasets. ScribblePrompt is more accurate than existing baselines, and ScribblePrompt-UNet is computationally efficient, even on a CPU. Our user study shows that nearly all users prefer ScribblePrompt and achieve segmentations with 15% higher Dice with less effort than the next most accurate baseline. ScribblePrompt promises to significantly reduce the burden of manual segmentation in biomedical imaging.
@@ 8 Acknowledgements
This work was supported in part by funding from Quanta Computer Inc., the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard and the Wistron Corporation. Research reported in this paper was supported by the National Institute of Biomedical Imaging and Bioengineering of the National Institutes of Health under award number R01EB033773. Much of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center.
@@
Appendix A: Demo and Code
Appendix B: ScribblePrompt Implementation
Appendix D: Experimental Setup
Appendix C: Data
Appendix E: Evaluation with Manual Scribbles
Appendix F: Evaluation with Simulated Interactions
Appendix G: User Study
Appendix H: Inference Runtime
Appendix I: Ablations
Appendix A Demo and Code
An interactive demo, code, model weights, and the MedScribble dataset are available at
https://scribbleprompt.csail.mit.edu
Appendix B ScribblePrompt Implementation
@@ B.1.1 Scribbles
Line Scribbles. Fig. 8 illustrates the process of simulating line scribbles.
Figure 8: Line scribbles. Given an input mask $z$, we draw random lines by sampling two end points from $\{(u,v)|z_{uv}=1\}$. We use a random deformation field to warp the line scribbles and then multiply by the binary input mask $z$ to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the ground truth label $y$ (top) and negative scribbles by applying the algorithm to the background $1-y$ (bottom).
Centerline Scribbles. Fig. 9 illustrates the process of simulating centerline scribbles.
@@
Contour Scribbles. Fig. 10 illustrates the process of simulating contour scribbles.
Figure 10: Contour scribbles. We simulate a rough contour of the desired segmentation within the boundaries of the label. Given a mask $z$, We first blur the mask to reduce the size of the label such that $\tilde{z}=\min(z,z\circ G_{k})$, where $G_{k}$ is a Gaussian blur kernel. Then we apply a threshold $\tilde{z}<h$ sampled in some intensity range $h\sim U[\tilde{z}_{min},\tilde{z}_{max}]$ and extract a contour inside the boundary of the mask. We break up the contour using a random mask and use a random deformation field to warp the broken contour. Lastly, we multiply the scribble mask by the input binary mask to correct parts of the scribble that were warped outside the mask. We can simulate positive scribbles by applying the algorithm to the label $y$ (bottom) and negative scribbles by applying the algorithm to the background $1-y$ (top).
@@ B.1.2 Clicks
Interior Border Region Clicks. Fig. 11 illustrates the process for simulating interior border region clicks.
Figure 11: Interior border region clicks. We sample clicks from a border region inside the boundary of a given mask. Given a mask $z$, we first blur the mask to reduce the size of the label such that $\tilde{z}=\min(z,z\circ G_{k})$ where $G_{k}$ is a Gaussian blur kernel. We then sample click coordinates from $\{(u,v)|\tilde{z}_{uv}\in[a,b]\}$, where $a,b\sim U[\tilde{z}_{min},\tilde{z}_{max})$ are thresholds sampled in some intensity range. We show the simulation process for negative border region clicks on the background $1-y$ (top) and positive border region clicks on the label $y$ (bottom).
@@ B.2 Architecture and Training
Normalization Layers. In preliminary experiments, we evaluated normalization layers in the ScribblePrompt-UNet architecture such as Batch Norm [54], Instance Norm [127], Layer Norm [8], and Channel Norm [139]. Including normalization did not improve the mean Dice on validation data compared to using no normalization layers (Fig. 12).
@@
Figure 12: Training ScribblePrompt-UNet with different normalization layers. We show mean Dice averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
@@
Loss Function. In preliminary experiments, we trained ScribblePrompt with Soft Dice Loss [32], a combination of Soft Dice Loss and Binary Cross-Entropy Loss, and a combination of Soft Dice Loss and Focal Loss [74], similar to [63]. In the latter two losses, Dice Loss and BCE Loss or Focal Loss are weighted equally. We found that the combination of Soft Dice Loss and Focal Loss resulted in slightly higher mean Dice on the validation data for ScribblePrompt-UNet and ScribblePrompt-SAM. Fig. 13 shows Dice recorded during training in preliminary experiments with ScribblePrompt-UNet.
@@
Figure 13: Training ScribblePrompt-UNet with different loss functions. We report Dice averaged across five iterative predictions (using the training procedure for simulating interactions). At each epoch, we evaluate on 1,000 randomly sampled examples from the validation splits of the 65 training datasets and validation splits of the nine validation datasets. Dice was smoothed using Exponential Weighted Mean with $\alpha=0.1$.
ScribblePrompt-UNet Inputs. We encode each prompt type in an input channel for ScribblePrompt-UNet. The input to ScribblePrompt-UNet has size $5\times h\times w$ consisting of the input image $x^{t}$, bounding box encoding, positive click/scribble encoding, negative click/scribble encoding, and the logits of the previous prediction $\hat{y}^{t}_{i-1}$. For the first prediction, we set the previous prediction channel to zeros. We encode bounding boxes in a binary mask that is 1 inside the box(es) and 0 everywhere else. We encode positive and negative clicks using binary masks where a pixel is 1 if has been clicked and 0 otherwise. We encode positive and negative scribbles as masks on $[0,1]$ and combine them with the masks encoding clicks. Representing the interactions as masks is advantageous because inference time does not scale with the number of interactions.
@@
ScribblePrompt-SAM Details. To train ScribblePrompt-SAM, we took the pre-trained weights from SAM [63] with ViT-b backbone and froze all components of the network except for the decoder.
@@
The SAM architecture can make predictions in single-mask mode or multi-mask mode. In single-mask mode, the decoder outputs a single predicted segmentation given an input image and user interactions. In multi-mask mode, the decoder predicts three possible segmentations and then outputs the segmentation with the highest predicted IoU by a MLP. We trained and evaluated ScribblePrompt-SAM in multi-mask mode to maximize the expressiveness of the architecture. During training we included a MSE term in the segmentation loss to train the MLP to predict the IoU of the predictions, as in [63].
@@ B.3 Synthetic Labels
To help reduce task overfitting – memorizing the segmentation task for single-label datasets and thus ignoring user prompts – we introduce a mechanism to generate synthetic labels. During training, for a given sample $(x_{0},y_{0})$, with probability $p_{synth}$ we replace $y_{0}$ with a synthetic label $y_{synth}$.
@@
We use a superpixel algorithm [33] with randomly sampled scale parameter $\lambda\sim U[1,500]$ to partition the image $x_{0}$ into a map of $k$ superpixels, $z\in\{1,\dots,k\}^{n\times n}$. Then, we randomly select a superpixel $c\sim\text{Cat}(\{1,\dots,k\},1/k)$ as the synthetic label $y_{synth}:=\mathbbm{1}[z=c]$. Fig. 14 shows examples of training images and the corresponding maps of possible synthetic labels with different $\lambda$.